OpenStudy (anonymous):

Which relationship shows an inverse variation? Choices: x = 1, 2, 3, 4 y = 3, 6, 9, 12 ------------- x = 1, 2, 3, 4 y = 60, 30, 20, 15 --------------- x = 1, 2, 3, 4 y = 15, 13, 11, 9 --------------- x = 1, 2, 3, 4 y = 6, 11, 20, 35

hartnn (hartnn):

find ,xy in each case, the one in which xy has same value, will have inverse variation.

OpenStudy (anonymous):

Can you give me an example?

hartnn (hartnn):

i'll do first choice for you. xy =1*3 =3 xy=2*6=12 not same , so its not inverse var.

hartnn (hartnn):

one more, last option : xy=1*6 =6 xy =2*11 = 22 not same, so not inverse variation.

hartnn (hartnn):

you try for other 2.

OpenStudy (anonymous):

x = 1, 2, 3, 4 y = 60, 30, 20, 15 I htink its this one

hartnn (hartnn):

yes! and the product everytime , xy =.... ?

OpenStudy (anonymous):

sorry my computer is going very slow its 5?Im just guessing here

hartnn (hartnn):

no... see, xy = 1*60 =60 xy = 2*30=60 xy=3*20=60 xy = 4*15=60 so, everytime xy=60 hence your function is xy=60 and is an inverse variation . if you were asked direct variation, you would have checked x/y=...

hartnn (hartnn):

understood ?